OpenStudy (anonymous):

PLEASE I am sure I got the equation. but not sure about the $16 plays in there. Jerome is contemplating the purchase of 100 shares of a stock selling $16 per share. The stock pays no dividends. The history of the stock indicates that it should grow at an annual rate of 12% per year. Use annual compounding to determine how much the 100 shares of stock will be worth in 8 years. The 100 shares will be worth $____ in 8 years

OpenStudy (anonymous):

\[100\left( 1+\frac{ 0.12 }{ 365 } \right)^{365\left( 8 \right)} \] I am not sure how the $16 plays in the equation.

OpenStudy (espex):

Generally speaking, it doesn't play a part. However, if you were to define "worth" as the amount of profit after 8 years, you would need to subtract the initial cost of the stock from the total value.

OpenStudy (anonymous):

so i got the answer 216.12 from the equation so do I subtract 100? and I will get 161.128

OpenStudy (espex):

The original purchase price was 100 shares @ 16 dollars each, giving you an initial expense of $1600. So rather than taking your initial investment as 100, you need to use the cost of the stocks.

OpenStudy (anonymous):

how would you do that??

OpenStudy (anonymous):

so your saying do 1600 - 216.12?

OpenStudy (espex):

I'm suggesting that instead of \[100\left( 1+\frac{ 0.12 }{ 365 } \right)^{365\left( 8 \right)}\] that your equation should be \[1600\left( 1+\frac{ 0.12 }{ 365 } \right)^{365\left( 8 \right)}\]

OpenStudy (espex):

With a compound interest formula of: \[A=P(1+\frac{r}{n})^{nt}\] paying once a year, for 8 years, I believe your equation should actually read: \[1600\left( 1+\frac{ 0.12 }{ 1 } \right)^{1\left( 8 \right)}\] and give you a total of 3961.54